# Li Tana,b , Shengrong Wanga,c

### v1

@@ 3 Existence and Uniqueness
$\left(\mathbb{W}_{q}(\mathcal{L}_{X_{r-1/n}^{n}},\delta_{0})\right)^{\overline{p}}=\left(\mathcal{W}_{q}(\mathcal{L}_{X_{r-1/n}^{n}})\right)^{\overline{p}}\leq\mathbb{E}|X_{r-1/n}^{n}|^{\overline{p}}.$
To sum up, we obtain
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|X_{t}^{n}-D(X_{t-\tau}^{n})|^{\overline{p}}\bigg{)}\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u-\tau}^{n}|^{(l+1)\overline{p}}\bigg{)}{\rm d}s,$
together with (3.2), we get
$\mathbb{E}\bigg{(}\sup_{0\leq r\leq t}|X_{r}^{n}|^{\overline{p}}\bigg{)}\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u-\tau}^{n}|^{(l+1)\overline{p}}\bigg{)}{\rm d}s.$
We set up a finite sequence $\{u_{1},u_{2},\cdots,u_{[T/\tau]+1}\}$ where $u_{i}=([T/\tau]+2-i)\overline{p}(l+1)^{[T/\tau]+1-i}$, and $[T/\tau]$ denotes the integer part of $T/\tau$. It is easy to see that $u_{i}\geq\overline{p}$, $u_{[T/\tau]+1}=\overline{p}$ and $(l+1)u_{i+1}<u_{i}$ for $i=1,2,\cdots,[T/\tau]$. For $s\in[0,\tau]$, we immediately get
$\mathbb{E}\bigg{(}\sup_{0\leq r\leq t}|X_{r}^{n}|^{\overline{p}}\bigg{)}\leq C,$
since $\xi\in L_{\mathcal{F}_{0}}^{\overline{p}}([-\tau,0];\mathbb{R}^{d})$. For $s\in[0,2\tau]$, by the Hölder inequality, we get

### v2

@@ 3 Existence and Uniqueness
(3.5) $\mathbb{E}\left(\sup_{0\leq s\leq t}|\Delta_{s}^{(k+1),(k)}|^{\overline{p}}\right)\leq\frac{1}{(1-\lambda)^{\overline{p}}}\mathbb{E}\left(\sup_{0\leq s\leq t}|\Upsilon_{s}^{(k+1),(k)}|^{\overline{p}}\right).$
For $t\in[0,T]$, employment of the fractional Itô formula in Lemma 2.1 and the elementary inequality provides
$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|\Upsilon_{s}^{(k+1),(k)}|^{\overline{p}}\bigg)$
$\leq$ ${\overline{p}}\mathbb{E}\int_{0}^{t}|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}-2}\langle\Upsilon_{r}^{(k+1),(k)},b(X^{(k+1)}_{r},X^{(k+1)}_{r-\tau},\mathcal{L}_{X^{(k)}_{r}})-b(X^{(k)}_{r},X^{(k)}_{r-\tau},\mathcal{L}_{X^{(k-1)}_{r}})\rangle{\rm d}r$
$+\mathbb{E}\bigg(\sup_{0\leq s\leq t}\bigg\lvert\int_{0}^{s}{\overline{p}}({\overline{p}}-1)H(2H-1)|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}-2}$
$\times\bigg(\int_{0}^{r}(r-u)^{2H-2}\|\sigma(\mathcal{L}_{X^{(k)}_{u}})-\sigma(\mathcal{L}_{X^{(k-1)}_{u}})\|^{2}{\rm d}u\bigg){\rm d}r\bigg)$
$+{\overline{p}}\mathbb{E}\bigg(\sup_{0\leq s\leq t}\int_{0}^{s}|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}-1}\|\sigma(\mathcal{L}_{X^{(k)}_{r}})-\sigma(\mathcal{L}_{X^{(k-1)}_{r}})\|{\rm d}B_{r}^{H}\bigg).$
$:=$ $J_{1}(t)+J_{2}(t)+J_{3}(t).$
Applying Assumption 2.4 and the Young’s inequality produces
$J_{1}(t)=$ ${\overline{p}}\mathbb{E}\int_{0}^{t}|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}-2}\langle\Upsilon_{r}^{(k+1),(k)},b(X^{(k+1)}_{r},X^{(k+1)}_{r-\tau},\mathcal{L}_{X^{(k)}_{r}})-b(X^{(k)}_{r},X^{(k)}_{r-\tau},\mathcal{L}_{X^{(k-1)}_{r}})\rangle{\rm d}r$
$\leq$ $C\mathbb{E}\int_{0}^{t}|X^{(k+1)}_{r}-D(X^{(k+1)}_{r-\tau})+X^{(k)}_{r}-D(X^{(k)}_{r-\tau})|^{\overline{p}}{\rm d}r$
$+C\mathbb{E}\int_{0}^{t}\bigg[|X^{(k+1)}_{r}-X^{(k)}_{r}|^{2}+|X^{(k+1)}_{r-\tau}-X^{(k)}_{r-\tau}|^{2}+\mathbb{W}_{2}^{2}(\mathcal{L}_{X^{(k)}_{r}},\mathcal{L}_{X^{(k-1)}_{r}})\bigg]^{{\overline{p}}/2}{\rm d}r$
$\leq$ $C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k+1),(k)}|^{\overline{p}}\bigg){\rm d}r+C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{\overline{p}}\bigg){\rm d}r,$
where we used the fact that $\mathbb{W}_{2}^{2}(\mathcal{L}_{X^{(k)}_{r}},\mathcal{L}_{X^{(k-1)}_{r}})\leq\mathbb{E}^{\mathbb{P}}|X^{(k)}_{r}-X^{(k-1)}_{r}|^{2}$. Along with Assumption 2.5, the Young’s inequality, the Hölder inequality, and the Fubini theorem, it follows
$J_{2}(t)=$ $\mathbb{E}\bigg(\sup_{0\leq s\leq t}\bigg\lvert\int_{0}^{s}{\overline{p}}({\overline{p}}-1)H(2H-1)|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}-2}$
$\times\left(\int_{0}^{r}(r-u)^{2H-2}\|\sigma(\mathcal{L}_{X^{(k)}_{u}})-\sigma(\mathcal{L}_{X^{(k-1)}_{u}})\|^{2}{\rm d}u\bigg){\rm d}r\right)$
$\leq$ $C\int_{0}^{t}\int_{0}^{r}(r-u)^{2H-2}(\mathbb{E}|\Upsilon_{r}^{(k+1),(k)}|^{\overline{p}}+\mathbb{E}|X^{(k)}_{u}-X^{(k-1)}_{u}|^{\overline{p}}){\rm d}u{\rm d}r$
$\leq$ $C\int_{0}^{t}\mathbb{E}|\Upsilon_{r}^{(k+1),(k)}|^{\overline{p}}\bigg(\int_{0}^{r}(r-u)^{2H-2}{\rm d}u\bigg){\rm d}r+C\int_{0}^{t}\int_{u}^{t}(r-u)^{2H-2}\mathbb{E}|X^{(k)}_{u}-X^{(k-1)}_{u}|^{\overline{p}}{\rm d}r{\rm d}u$
$\leq$ $C\int_{0}^{t}\frac{r^{2H-1}}{2H-1}\mathbb{E}|\Upsilon_{r}^{(k+1),(k)}|^{\overline{p}}{\rm d}r+C\int_{0}^{t}\frac{(t-u)^{2H-1}}{2H-1}\mathbb{E}|X^{(k)}_{u}-X^{(k-1)}_{u}|^{\overline{p}}{\rm d}u.$
Since $1/2<H<1$ and $t\in[0,T]$, we use Assumption 2.3 to get
$J_{2}(t)\leq$ $C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k+1),(k)}|^{\overline{p}}\bigg){\rm d}r+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{\overline{p}}\bigg){\rm d}r.$
In the similar way as dealing with $Q_{3}^{(1)}(t)$ to get (3.4), an application of fractional BDG inequality (3.3) with $p=1$ yields
$J_{3}(t)=$ ${\overline{p}}\mathbb{E}\left[\sup_{0\leq s\leq t}\left|\int_{0}^{s}|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}-1}\|\sigma(\mathcal{L}_{X^{(k)}_{r}})-\sigma(\mathcal{L}_{X^{(k-1)}_{r}})\|{\rm d}B_{r}^{H}\right|\right]$
$\leq$ $C\mathbb{E}\left[\left(\int_{0}^{t}\left(|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}-1}\|\sigma(\mathcal{L}_{X^{(k)}_{r}})-\sigma(\mathcal{L}_{X^{(k-1)}_{r}})\|\right)^{1/H}{\rm d}r\right)^{H}\right]$
For each $r\in[0,t]$, by the Young’s inequality, there exists $0<\epsilon<1$ such that
$|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}-1}\|\sigma(\mathcal{L}_{X^{(k)}_{r}})-\sigma(\mathcal{L}_{X^{(k-1)}_{r}})\|\leq\epsilon|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}}+C_{\epsilon}\|\sigma(\mathcal{L}_{X^{(k)}_{r}})-\sigma(\mathcal{L}_{X^{(k-1)}_{r}})\|^{{\overline{p}}}.$
Assumption 2.5 implies that
$\|\sigma(\mathcal{L}_{X^{(k)}_{r}})-\sigma(\mathcal{L}_{X^{(k-1)}_{r}})\|\leq K_{4}\mathbb{W}_{2}(\mathcal{L}_{X^{(k)}_{r}},\mathcal{L}_{X^{(k-1)}_{r}})\leq K_{4}(\mathbb{E}|\Delta_{r}^{(k),(k-1)}|^{2})^{1/2}.$
Since ${\overline{p}}\geq 2$, using the fact that $(\mathbb{E}|\Upsilon_{r}^{(k),(k-1)}|^{2})^{{\overline{p}}/2}\leq\mathbb{E}|\Upsilon_{r}^{(k),(k-1)}|^{{\overline{p}}}$, then
$|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}-1}\|\sigma(\mathcal{L}_{X^{(k)}_{r}})-\sigma(\mathcal{L}_{X^{(k-1)}_{r}})\|\leq\epsilon|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}}+C_{\epsilon}(\mathbb{E}|\Delta_{r}^{(k),(k-1)}|^{2})^{{\overline{p}}/2}$
$\leq$ $\epsilon|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}}+C_{\epsilon}\mathbb{E}|\Delta_{r}^{(k),(k-1)}|^{{\overline{p}}}.$
Noting that for $\frac{1}{H}>1$, we can apply the elementary inequality to get
$\left(\int_{0}^{t}\left(\epsilon|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}}+C_{\epsilon}\mathbb{E}|\Delta_{r}^{(k),(k-1)}|^{{\overline{p}}}\right)^{1/H}{\rm d}r\right)^{H}$
$\leq$ $2^{1-H}\epsilon\left(\int_{0}^{t}|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}/H}{\rm d}r\right)^{H}+C_{\epsilon}2^{1-H}\left(\int_{0}^{t}\left(\mathbb{E}|\Delta_{r}^{(k),(k-1)}|^{{\overline{p}}}\right)^{1/H}{\rm d}r\right)^{H}.$
Taking expectation, it holds
$\mathbb{E}\left[\left(\int_{0}^{t}|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}/H}{\rm d}r\right)^{H}\right]\leq t^{H}\mathbb{E}\left(\sup_{0\leq r\leq t}|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}}\right),$
and
$\mathbb{E}\left[\left(\int_{0}^{t}\left(\mathbb{E}|\Delta_{r}^{(k),(k-1)}|^{{\overline{p}}}\right)^{1/H}{\rm d}r\right)^{H}\right]\leq Ct^{H}\int_{0}^{t}\mathbb{E}|\Delta_{r}^{(k),(k-1)}|^{{\overline{p}}}{\rm d}r.$
Combining the estimates together, we conclude
(3.6) $J_{3}(t)\leq$ $C\epsilon\mathbb{E}\left(\sup_{0\leq s\leq t}|\Upsilon_{s}^{(k+1),(k)}|^{{\overline{p}}}\right)+C_{\epsilon}\int_{0}^{t}\mathbb{E}\left(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{{\overline{p}}}\right){\rm d}r.$
Finally, the estimation of $J_{3}(t)$ combined with the estimates for the drift terms $J_{1}(t)$ and $J_{2}(t)$, we obtain
$\mathbb{E}\left(\sup_{0\leq s\leq t}|\Upsilon_{s}^{(k+1),(k)}|^{\overline{p}}\right)$